OpenStudy (aundiewear):

The expression tan^3(θ) + (sin(θ)/cos(θ)) is equal to: a) cot(θ) b) tan(θ)sec^2(θ) c) tan(θ) d) sin(θ)tan(θ) e) tan(θ)csc^2(θ)

OpenStudy (kittiwitti1):

\[\tan^{3}{\theta}+(\frac{\sin{\theta}}{\cos{\theta}})=?\]

OpenStudy (kittiwitti1):

\[\text{Rule: }\frac{\sin{\theta}}{\cos{\theta}}=\tan{\theta}\]\[\text{SO we get }\tan^{3}{\theta}+\tan{\theta}...\]

OpenStudy (kittiwitti1):

\[\text{This leads to: }\tan{\theta}(\tan^{2}{\theta}+1)\]

OpenStudy (kittiwitti1):

\[\text{What identity is }"\tan^{2}{\theta}+1"?\]

OpenStudy (kittiwitti1):

http://www.sosmath.com/trig/Trig5/trig5/trig5.html

OpenStudy (aundiewear):

sec^2(θ)

OpenStudy (aundiewear):

thank you so much

OpenStudy (kittiwitti1):

Don't forget the tangent!\[A:\tan{\theta}\sec^{2}{\theta}\]